OpenStudy (anonymous):

The pH factor determines whether a substance is alkaline, neutral or acidic depending if pH is greater than, equal to or less than 7 respectively. If H+ is the hydrogen ion concentration in moles, pH is determined by the formula pH = − log(H + ) . For tomato juice, H+ = 6.3 x 10^-5 moles. In which of the following intervals does pH fall for tomato juice? A. [1,2] B. [4,5] C. [5,6] D. [11,12] E. NOTA Please Explain! Also, No Calculator!

OpenStudy (anonymous):

the ans would be B.{4,5}

OpenStudy (anonymous):

Why?

OpenStudy (anonymous):

ph=-log(+h) =-log(6.3x10^-5) =-[log(6.3 ) + log(10^-5)] =-[0.799+ (-5) ] = -[-4.2] =4.2

OpenStudy (anonymous):

I can't use a calculator for this though.

OpenStudy (foolaroundmath):

ph = -log([H+]) = -log(6.3*10^(-5)) = -[log(6.3) + log(10^-5)] = -log(6.3) -(-5) = 5 - log(6.3) Now 1 < 6.3 < 10 so, log1 < log6.3 < log10 => 0 < log6.3 < 1 thus pH = 5 - x and x is between 0 and 1 Thus, pH lies in between 4,5

OpenStudy (anonymous):

Oh ok. I knew I was close, but I couldn't quite finish it. THanks @FoolAroundMath !